OpenStudy (anonymous):

solve for n. (6 1/2 + 5 1/5) + 2 4/5= n + 10 I REALLY NEED LIKE IMMEDIATE HELP!!!!!!!!!!

OpenStudy (tommynaut):

Your equation is a little bit unclear - did you mean this? \[\left( 6\frac{ 1 }{ 2 } + 5 \frac{ 1 }{ 5 } \right) + 2\frac{ 4 }{ 5 } = n +10\]

OpenStudy (tommynaut):

Or did you mean "6 times a half plus 5 times a fifth... etc"?

OpenStudy (triciaal):

one approach separate the whole numbers and the fractions to add separately then add the 2 sub-totals

OpenStudy (tommynaut):

If you meant the first scenario, to solve for n you just have to subtract 10 from both sides of the equation. That would get rid of the 10 from the right hand side so you only have n on that side, but what you do to one side of an equation you have to do to the other, so you would have an additional -10 on the left hand side.

OpenStudy (triciaal):

you have an equation to solve for the variable isolate and whatever you do to one side also do to the other to maintain the balance

OpenStudy (triciaal):

what did you get for n ? do you want me to check?